what % of 125 is 17
Answer by stanbon(75887)

what % of 125 is 17
---
x/100 = 17/125
---
x = 100(17/125)
---
x = (4/5)17
---
x = 13.6
======
Ans 13.6%
==============
Cheers,
Stan H.